# Cointegration with Occasionally Binding Constraints

### v3

@@ Appendix B Proofs of theorems
(i). Recall from (LABEL:pikink) that $\beta^{\pm\mathsf{T}}=[\beta_{y}^{\pm},\beta_{x}^{\mathsf{T}}]=\beta_{x}^{\mathsf{T}}[\theta^{\pm},I_{p-1}]$ for some $\theta^{\pm}\in\mathbb{R}^{p-1}$, and take
@@
as per (LABEL:betaperpfn), so that $e_{1}^{\mathsf{T}}\beta_{\perp}^{\pm}=e_{1}^{\mathsf{T}}$; it follows that $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)\neq 0$. Since
@@
where $\delta_{\gamma}\coloneqq[\gamma^{-}(1)-\gamma^{+}(1)-\Gamma^{x}(1)(\theta^{-}-\theta^{+})]$. Under CO(ii).3, we may apply A.2 with $A=\alpha_{\perp}^{\mathsf{T}}$, $B_{1}=\Gamma^{-}(1)\beta_{\perp}^{-}$, $B_{2}=\Gamma^{+}(1)\beta_{\perp}^{+}$ and $d=e_{1}$, to obtain
@@
where $v\coloneqq\alpha_{\perp}^{\mathsf{T}}u$, and $A$, $B_{1}$, $B_{2}$ and $d$ are as above. Hence by A.2,
00 0 $=[(A^{\mathsf{T}}B_{2})^{-1}-(A^{\mathsf{T}}B_{1})^{-1}]v=[(\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+})^{-1}-(\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-})^{-1}]\alpha_{\perp}^{\mathsf{T}}u.$
Finally, noting that (LABEL:betaperp) implies $\beta_{\perp}^{+}-\beta_{\perp}^{-}=\delta_{\beta}e_{1}^{\mathsf{T}}$ for $\delta_{\beta}\coloneqq(0,(\theta^{-}-\theta^{+})^{\mathsf{T}})^{\mathsf{T}}$, we have
@@
where $=_{(1)}$ holds by $e_{1}^{\mathsf{T}}\beta_{\perp}^{-}=e_{1}^{\mathsf{T}}$, $=_{(2)}$ by the result of part (i), and $=_{(3)}$ by our choice of $u$.
@@
where $C_{0}\coloneqq\max\{1,\mu^{-1}\}\in(0,\infty)$. (If (b) holds instead, then (LABEL:vartheta-u) follows by an analogous argument.) Now let $u_{\vartheta_{\perp}}\coloneqq u-\vartheta(\vartheta^{\mathsf{T}}\vartheta)^{-1}\vartheta^{\mathsf{T}}u$. Then $\vartheta^{\mathsf{T}}u_{\vartheta_{\perp}}=0$, and so
@@
where the second equality follows by the result of part (ii). By (LABEL:vartheta-u),
@@
Thus (LABEL:gLip) follows from (LABEL:gdiff)–(LABEL:uprime-utheta) ∎
Returning now to (LABEL:beforewkc), we note that (LABEL:detprime-case2) and B.3 imply that $P_{\beta_{\perp}}(y)c=P_{\beta_{\perp}}(+1)c$ for all $y\in\mathbb{R}$. Therefore by defining the detrended process
@@
for $E_{-1}$ the final $p-1$ columns of $I_{p}$, we may rewrite (LABEL:beforewkc) as
@@
It remains to determine the weak limit of the r.h.s. of (LABEL:wkcoclas). By B.3, $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(y)=h(y)\vartheta^{\mathsf{T}}$, and the first equation of (LABEL:wkcoclas) becomes
@@
on $D[0,1]$. Now consider the remaining $p-1$ equations in (LABEL:wkcoclas). Recalling the definition of $g$ in (LABEL:gdef) above, we have
@@
By (LABEL:Ylim), $(Y_{n},U_{n,0})\rightsquigarrow(Y,U_{0})$, which have continuous paths and concentrate in the set
@@
By B.3, if $(y_{n},u_{n})\in D_{\mathbb{R}^{p+1}}[0,1]$ converge uniformly to $(y,u)\in\overline{D}$, then $g[y_{n}(\lambda),u_{n}(\lambda)]\rightarrow g[y(\lambda),u(\lambda)]$ uniformly over $\lambda\in[0,1]$. Hence by the CMT,
@@
Finally, we suppose that DET holds in place of (LABEL:detprime-case2). In this case, $P_{\beta_{\perp}}(+1)c=0$ and therefore $x_{t}^{d}=x_{t}$ for all $t\in\{1,\ldots,T\}$. Thus part (ii) of the theorem follows from (LABEL:Ylim) and (LABEL:Xlim). ∎
B.3 Proof of 4.3
@@
To prove 4.3, we need to put the model into a VECM form different again from that used in the analysis of the preceding cases. Recalling $z_{t}^{\mathsf{T}}=(y_{t},x_{t}^{\mathsf{T}})$, the canonical CKSVAR (LABEL:canon-var) may be written as
@@
where $\Pi(\boldsymbol{y}_{t-i})\coloneqq\sum_{i=1}^{k}\Phi_{i}(y_{t-i})-I_{p}$, and as per (LABEL:gamstatinitial) above,
@@
where for $\boldsymbol{\alpha}(\boldsymbol{y}_{t-1})$ and $\boldsymbol{\beta}$ as defined in (LABEL:alphbetcase3),
@@
Hence for $\boldsymbol{\xi}_{t}\coloneqq\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t}=(\xi_{t}^{\mathsf{T}},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-k+2}^{\mathsf{T}})^{\mathsf{T}}$, where $\xi_{t}=\beta^{\mathsf{T}}z_{t}$, it follows from (LABEL:compcase3) that
@@
To ‘close’ (LABEL:bxistat), we need only to recognise that with the ordering of the cointegrating vectors given in (LABEL:costatfactor), $y_{t-1}=e_{r,1}^{\mathsf{T}}\xi_{t-1}$ and
@@
We would like to apply B.4 to this system, but we are prevented from doing so because $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}$ does not satisfy the required condition on the innovations. To remedy this, we modify the state vector of the system as follows. Let $\overline{\beta}\coloneqq\beta(\beta^{\mathsf{T}}\beta)^{-1}$, $\overline{\beta}_{\perp}\coloneqq\beta_{\perp}(\beta_{\perp}^{\mathsf{T}}\beta_{\perp})^{-1}$, and recall $\chi_{t}=(\xi_{t}^{\mathsf{T}},(\beta_{\perp}^{\mathsf{T}}\Delta z_{t})^{\mathsf{T}})^{\mathsf{T}}$ and $\boldsymbol{\chi}_{t}=(\chi_{t}^{\mathsf{T}},\ldots,\chi_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$. Since the r.h.s. of
@@
we have $M\boldsymbol{\xi}_{t}=\chi_{t}$, $M\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{c}=Bc$ and $M\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}=Bu_{t}$, and so premultiplying (LABEL:xisystem) by $M$ yields
@@
We may now apply B.4 to (LABEL:chisystem). Since $B$ has full rank, ERR is satisfied with $Bu_{t}$ taking the place of $u_{t}$. Regarding the other conditions of the lemma, let $\boldsymbol{F}$ be such that $\boldsymbol{F}\boldsymbol{z}_{t-1}=\boldsymbol{y}_{t-1}$. By the continuity of the r.h.s. of the CKSVAR, the map
@@
which simply replicates the autoregressive part of the companion form (LABEL:compcase3) of the CKSVAR (in levels), is continuous. Hence so too is
@@
$\boldsymbol{\xi}_{t-1}=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t-1}$ varies freely, and by our arguments following (LABEL:xisystem), $\boldsymbol{F}\boldsymbol{z}_{t-1}$ depends (continuously) only on the elements of $\boldsymbol{\xi}_{t-1}$. Indeed, $\boldsymbol{F}\boldsymbol{z}_{t-1}=\boldsymbol{y}_{t-1}=\boldsymbol{G}_{\xi}\boldsymbol{\xi}_{t-1}$ where $\boldsymbol{G}_{\xi}$ is as defined in (LABEL:xitoy). Hence the mapping
@@
that appears on the r.h.s. of (LABEL:xisystem), is also continuous. Thus
@@
is continuous by the composition of continuous maps. Since $\boldsymbol{\alpha}(\boldsymbol{y})$ depends only on the signs of the elements of $\boldsymbol{y}$, the preceding is also homogeneous of degree one. Finally, since it is possible to write $\boldsymbol{\chi}_{t+1}$ as a (deterministic) function of $\xi_{t+1}$ and $\boldsymbol{\xi}_{t}$, the deterministic system associated to (LABEL:chisystem) must be stable if that associated to (LABEL:xisystem) is stable, where the latter is maintained under CO(iii).2. The result now follows by B.4. ∎
B.4 Proof of 4.4
Suppose initially that DGP holds. With the exception of the final claim made in each, the proofs of Theorems 4.1 and 4.2 proceed under conditions (LABEL:detprime-case1) and (LABEL:detprime-case2) respectively, which correspond to DET rather than DET, and yield the claimed limit weak limits for $X_{n}^{d}$. Thus for $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$ generated by a canonical CKSVAR, the theorem is proved.
Now suppose that merely DGP holds (along with CO(i) or CO(ii)). It is shown in D that, in these cases, the derived canonical form (denoted by tildes) satisfies $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}\tilde{c}=0$ in case (i), and $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}}(+1)\tilde{c}=0$ in case (ii), i.e. DET holds in the canonical form; the argument subsequently given in that appendix shows that the conclusions of the theorem may then be transposed back to the structural form. ∎
@@ Appendix C Proofs of auxiliary lemmas
Proof of A.1.
Since $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})<1$, it follows by Jungers (2009, Prop. 1.4) that there exists a norm $\lVert\cdot\rVert_{\ast}$ on $\mathbb{R}^{d_{w}}$ and a $\gamma\in(0,1)$ such that
@@
[…]
@@
$d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}$ =$=\left\{1-\frac{d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}{1+d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}\right\}d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}=\chi^{-1}d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}$
and thus part (i) holds with $\mu=\chi^{-1}$. Finally, let $v\in\mathbb{R}^{n}$ be such that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}v=0$. By part (i) of the lemma, we must have that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}v=0$; hence postmultiplying (LABEL:smw) by $v$ immediately yields the result of part (ii). ∎
Proof of A.3.
The factorisation of $\Pi$ is immediate from the definition of matrix rank; that of $\boldsymbol{\Pi}$ then follows from direct calculation. Parts (iii) and (iv) will also follow from direct calculation, once we have verified (ii).
It remains to prove part (ii). We first observe that $I_{kp}+\boldsymbol{\Pi}$ gives the companion form matrix associated with the autoregressive polynomial $\Phi(\lambda)$, in the sense that the roots of the latter coincide with the reciprocals of the eigenvalues of the former (e.g. Lütkepohl, 2007, Sec. 2.1.1), of which $q$ lie on the unit circle by assumption, and the remaining $kp-q$ lie strictly inside. It is readily verified that $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\alpha}=0$ and $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}=0$, when $\boldsymbol{\alpha}_{\perp}$ and $\boldsymbol{\beta}_{\perp}$ have the form given in (LABEL:abperp). Moreover, since $\beta_{\perp}$ is determined only up to its column span, we are free to normalise it such that $\beta_{\perp}^{\mathsf{T}}\beta_{\perp}=k^{-1}I_{q}$, whence $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=I_{q}$. Letting $Q\coloneqq(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1/2}$ denote the positive definite square root of $(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1}$, we have that $[\boldsymbol{\beta}Q,\boldsymbol{\beta}_{\perp}]$ is a (full rank) orthogonal matrix. Since
@@
Hence the eigenvalues of $I_{kp-q}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ are the $kp-q$ eigenvalues of $I_{kp}+\boldsymbol{\Pi}$ that are strictly inside the unit circle. Deduce $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ has full rank, whence so does $[\boldsymbol{\beta}_{\perp},\boldsymbol{\alpha}]$. Since
$\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\begin{bmatrix}\boldsymbol{\beta}_{\perp}&\boldsymbol{\alpha}\end{bmatrix}=\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}&0\end{bmatrix}$
must then have rank $q$, it follows that the $q\times q$ matrix $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ has full rank. ∎

### v4

@@ Appendix B Proofs of theorems
(i). Recall from (4.6) that $\beta^{\pm\mathsf{T}}=[\beta_{y}^{\pm},\beta_{x}^{\mathsf{T}}]=\beta_{x}^{\mathsf{T}}[\theta^{\pm},I_{p-1}]$ for some $\theta^{\pm}\in\mathbb{R}^{p-1}$, and take
@@
as per (4.8), so that $e_{1}^{\mathsf{T}}\beta_{\perp}^{\pm}=e_{1}^{\mathsf{T}}$; it follows that $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)\neq 0$. Since
@@
where $\delta_{\gamma}\coloneqq[\gamma^{-}(1)-\gamma^{+}(1)-\Gamma^{x}(1)(\theta^{-}-\theta^{+})]$. Under CO(ii).3, we may apply Lemma A.2 with $A=\alpha_{\perp}^{\mathsf{T}}$, $B_{1}=\Gamma^{-}(1)\beta_{\perp}^{-}$, $B_{2}=\Gamma^{+}(1)\beta_{\perp}^{+}$ and $d=e_{1}$, to obtain
@@
where $v\coloneqq\alpha_{\perp}^{\mathsf{T}}u$, and $A$, $B_{1}$, $B_{2}$ and $d$ are as above. Hence by Lemma A.2,
$0$ $=[(A^{\mathsf{T}}B_{2})^{-1}-(A^{\mathsf{T}}B_{1})^{-1}]v=[(\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+})^{-1}-(\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-})^{-1}]\alpha_{\perp}^{\mathsf{T}}u.$
Finally, noting that (B.26) implies $\beta_{\perp}^{+}-\beta_{\perp}^{-}=\delta_{\beta}e_{1}^{\mathsf{T}}$ for $\delta_{\beta}\coloneqq(0,(\theta^{-}-\theta^{+})^{\mathsf{T}})^{\mathsf{T}}$, we have
@@
where $_{(1)}$ holds by $e_{1}^{\mathsf{T}}\beta_{\perp}^{-}=e_{1}^{\mathsf{T}}$, $_{(2)}$ by the result of part (i), and $_{(3)}$ by our choice of $u$.
@@
where $C_{0}\coloneqq\max\{1,\mu^{-1}\}\in(0,\infty)$. (If (b) holds instead, then (B.27) follows by an analogous argument.) Now let $u_{\vartheta_{\perp}}\coloneqq u-\vartheta(\vartheta^{\mathsf{T}}\vartheta)^{-1}\vartheta^{\mathsf{T}}u$. Then $\vartheta^{\mathsf{T}}u_{\vartheta_{\perp}}=0$, and so
@@
where the second equality follows by the result of part (ii). By (B.27),
@@
Thus (B.25) follows from (B.28)–(B.30) ∎
Returning now to (B.23), we note that (B.20) and Lemma B.3 imply that $P_{\beta_{\perp}}(y)c=P_{\beta_{\perp}}(+1)c$ for all $y\in\mathbb{R}$. Therefore by defining the detrended process
@@
for $E_{-1}$ the final p1p-1 columns of $I_{p}$, we may rewrite (B.23) as
@@
It remains to determine the weak limit of the r.h.s. of (B.31). By Lemma B.3, $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(y)=h(y)\vartheta^{\mathsf{T}}$, and the first equation of (B.31) becomes
@@
on $D[0,1]$. Now consider the remaining p1p-1 equations in (B.31). Recalling the definition of $g$ in (B.24) above, we have
@@
By (B.32), $(Y_{n},U_{n,0})\rightsquigarrow(Y,U_{0})$, which have continuous paths and concentrate in the set
@@
By Lemma B.3, if $(y_{n},u_{n})\in D_{\mathbb{R}^{p+1}}[0,1]$ converge uniformly to $(y,u)\in\overline{D}$, then $g[y_{n}(\lambda),u_{n}(\lambda)]\rightarrow g[y(\lambda),u(\lambda)]$ uniformly over $\lambda\in[0,1]$. Hence by the CMT,
@@
Finally, we suppose that DET holds in place of (B.20). In this case, $P_{\beta_{\perp}}(+1)c=0$ and therefore $x_{t}^{d}=x_{t}$ for all $t\in\{1,\ldots,T\}$. Thus part (ii) of the theorem follows from (B.32) and (B.33). ∎
B.3 Proof of Theorem 4.3
@@
To prove Theorem 4.3, we need to put the model into a VECM form different again from that used in the analysis of the preceding cases. Recalling $z_{t}^{\mathsf{T}}=(y_{t},x_{t}^{\mathsf{T}})$, the canonical CKSVAR (2.17) may be written as
@@
where $\Pi(\boldsymbol{y}_{t-i})\coloneqq\sum_{i=1}^{k}\Phi_{i}(y_{t-i})-I_{p}$, and as per (4.17) above,
@@
where for $\boldsymbol{\alpha}(\boldsymbol{y}_{t-1})$ and $\boldsymbol{\beta}$ as defined in (4.16),
@@
Hence for $\boldsymbol{\xi}_{t}\coloneqq\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t}=(\xi_{t}^{\mathsf{T}},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-k+2}^{\mathsf{T}})^{\mathsf{T}}$, where $\xi_{t}=\beta^{\mathsf{T}}z_{t}$, it follows from (B.35) that
@@
To ‘close’ (B.36), we need only to recognise that with the ordering of the cointegrating vectors given in (4.15), $y_{t-1}=e_{r,1}^{\mathsf{T}}\xi_{t-1}$ and
@@
We would like to apply Lemma B.4 to this system, but we are prevented from doing so because $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}$ does not satisfy the required condition on the innovations. To remedy this, we modify the state vector of the system as follows. Let $\overline{\beta}\coloneqq\beta(\beta^{\mathsf{T}}\beta)^{-1}$, $\overline{\beta}_{\perp}\coloneqq\beta_{\perp}(\beta_{\perp}^{\mathsf{T}}\beta_{\perp})^{-1}$, and recall $\chi_{t}=(\xi_{t}^{\mathsf{T}},(\beta_{\perp}^{\mathsf{T}}\Delta z_{t})^{\mathsf{T}})^{\mathsf{T}}$ and $\boldsymbol{\chi}_{t}=(\chi_{t}^{\mathsf{T}},\ldots,\chi_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$. Since the r.h.s. of
@@
we have $M\boldsymbol{\xi}_{t}=\chi_{t}$, $M\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{c}=Bc$ and $M\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}=Bu_{t}$, and so premultiplying (B.38) by $M$ yields
@@
We may now apply Lemma B.4 to (B.39). Since $B$ has full rank, ERR is satisfied with $Bu_{t}$ taking the place of $u_{t}$. Regarding the other conditions of the lemma, let $\boldsymbol{F}$ be such that $\boldsymbol{F}\boldsymbol{z}_{t-1}=\boldsymbol{y}_{t-1}$. By the continuity of the r.h.s. of the CKSVAR, the map
@@
which simply replicates the autoregressive part of the companion form (B.35) of the CKSVAR (in levels), is continuous. Hence so too is
@@
$\boldsymbol{\xi}_{t-1}=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t-1}$ varies freely, and by our arguments following (B.38), $\boldsymbol{F}\boldsymbol{z}_{t-1}$ depends (continuously) only on the elements of $\boldsymbol{\xi}_{t-1}$. Indeed, $\boldsymbol{F}\boldsymbol{z}_{t-1}=\boldsymbol{y}_{t-1}=\boldsymbol{G}_{\xi}\boldsymbol{\xi}_{t-1}$ where $\boldsymbol{G}_{\xi}$ is as defined in (B.37). Hence the mapping
@@
that appears on the r.h.s. of (B.38), is also continuous. Thus
@@
is continuous by the composition of continuous maps. Since $\boldsymbol{\alpha}(\boldsymbol{y})$ depends only on the signs of the elements of $\boldsymbol{y}$, the preceding is also homogeneous of degree one. Finally, since it is possible to write $\boldsymbol{\chi}_{t+1}$ as a (deterministic) function of $\xi_{t+1}$ and $\boldsymbol{\xi}_{t}$, the deterministic system associated to (B.39) must be stable if that associated to (B.38) is stable, where the latter is maintained under CO(iii).2. The result now follows by Lemma B.4. ∎
B.4 Proof of Theorem 4.4
Suppose initially that DGP holds. With the exception of the final claim made in each, the proofs of Theorems 4.1 and 4.2 proceed under conditions (B.8) and (B.20) respectively, which correspond to DET rather than DET, and yield the claimed limit weak limits for $X_{n}^{d}$. Thus for $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$ generated by a canonical CKSVAR, the theorem is proved.
Now suppose that merely DGP holds (along with CO(i) or CO(ii)). It is shown in Appendix D that, in these cases, the derived canonical form (denoted by tildes) satisfies $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}\tilde{c}=0$ in case (i), and $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}}(+1)\tilde{c}=0$ in case (ii), i.e. DET holds in the canonical form; the argument subsequently given in that appendix shows that the conclusions of the theorem may then be transposed back to the structural form. ∎
@@ Appendix C Proofs of auxiliary lemmas
Proof of Lemma A.1.
Since $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})<1$, it follows by Jungers (2009, Prop. 1.4) that for any given $\gamma\in(\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A}),1)$, there exists a norm $\lVert\cdot\rVert_{\ast}$ on $\mathbb{R}^{d_{w}}$ such that
@@
where $\lVert\cdot\rVert$ denotes the Euclidean norm on $\mathbb{R}^{d_{w}}$. Deduce that for any $m\in[1,m_{0}]$,
@@
$d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}$ $=\left\{1-\frac{d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}{1+d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}\right\}d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}=\chi^{-1}d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}$
and thus part (i) holds with $\mu=\chi^{-1}$. Finally, let $v\in\mathbb{R}^{n}$ be such that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}v=0$. By part (i) of the lemma, we must have that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}v=0$; hence postmultiplying (C.1) by $v$ immediately yields the result of part (ii). ∎
Proof of Lemma A.3.
The factorisation of $\Pi$ is immediate from the definition of matrix rank; that of $\boldsymbol{\Pi}$ then follows from direct calculation. Part (iii) will also follow from direct calculation, once we have verified (ii).
It remains to prove part (ii), which is the converse to Lemma A.2 in Hansen (2005). We first observe that $I_{kp}+\boldsymbol{\Pi}$ gives the companion form matrix associated with the autoregressive polynomial $\Phi(\lambda)$, in the sense that the roots of the latter coincide with the reciprocals of the eigenvalues of the former (e.g. Lütkepohl, 2007, Sec. 2.1.1), of which $q$ lie on the unit circle by assumption, and the remaining kpqkp-q lie strictly inside. It is readily verified that $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\alpha}=0$ and $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}=0$, when $\boldsymbol{\alpha}_{\perp}$ and $\boldsymbol{\beta}_{\perp}$ have the form given in (A.1). Moreover, since $\beta_{\perp}$ is determined only up to its column span, we are free to normalise it such that $\beta_{\perp}^{\mathsf{T}}\beta_{\perp}=k^{-1}I_{q}$, whence $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=I_{q}$. Letting $Q\coloneqq(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1/2}$ denote the positive definite square root of $(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1}$, we have that $\boldsymbol{\beta}Q,\boldsymbol{\beta}_{\perp}]$ is a (full rank) orthogonal matrix. Since
@@
Hence the eigenvalues of $I_{kp-q}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ are the kpqkp-q eigenvalues of $I_{kp}+\boldsymbol{\Pi}$ that are strictly inside the unit circle. It then follows by Lemma A.1 in Hansen (2005) that $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ and $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ have full rank. ∎